Документ взят из кэша поисковой машины. Адрес оригинального документа : http://www.stsci.edu/smov/notes/0416smov.html
Дата изменения: Sat Apr 19 01:18:55 1997
Дата индексирования: Thu Feb 27 23:56:19 2014
Кодировка:

Поисковые слова: http www.badastronomy.com phpbb index.php
SMOV Status & Planning Meeting Minutes 04/16/97

SMOV Status & Planning Meeting Minutes

Meeting Date: 04/16/97


		SMOV STATUS & PLANNING MEETING -  16 APR. 97



NOTE:  The next meeting is April 18, 8:30 AM.


STATUS REPORTS

1.  Observatory Status - All systems nominal.

	DMS reported a problem last night in managing an SSR dump.  In
	preparing for an ED dump, the dump pointer was inadvertently placed
	behind the record pointer by 2 data blocks.  In attempting to
	recover from the misconfiguration, it was decided to transfer
	the ED to an ESTR, at which point a mis-directed command resulted
	in some science data loss.  The science data loss involved STIS
	7068 Diffuse Source Acq Test.  The exact amount and times of 
	the data loss were not known at the time of the meeting.


2.  WFPC2 Status - Proceeding nominally with Cycle 6 science and
	remaining SMOV monitors.  
  

3.  NICMOS Status

	- Engineering Status - Nominal
	The cold well (ice) temperature is 59.9 degK, stable.
	The VCS Cold Mask temperature is 97.7 degK., and appears stable.


  	- Science Status - 
	The 7043 Focus Monitor ran yesterday but the analysis is not yet
	completed.

	Data from the 7034 FOM test is now in hand for doing re-calibration
	to test FOM repeatability and PSFs (See notes from April 14)

	Camera 2 has displayed a increase in the thermal background along
	one edgs of the FOV.  It shows up as about a 20% increase in
	background brightness relative to the rest of the camera FOV.  This
	may be due to emission from of the Field Divider Mirror, which in
	turn may be caused by the detector's detected lateral shift.
	Investigation is continuing.
 
	7053 Limb Avoidance Proposal has completed but data receipt and
	analysis have not yet begun.
 
4.  STIS Status

	- Engineering Status - Nominal.  

	- Science Status - The impact of last night's data loss on
	7068 Diffuse Source Acq (See Observatory status above.) is not 			yet known.  Steve Kraemer will provide an assessment at Friday's
	meeting.

5.  FGS status - No new activities.

6.  FOC/COSTAR status- The FOC SMOV program is complete.  The SMOV TTRB has
	enabled the resumption of FOC science. 


PLANNING REPORTS

1.  Status of Upcoming SMSs


	- SMS104 - Currently executing.

	- SMS111 - Final Mode Products are out for review.

	- SMS118 - Needs to be rerun later in the week to include commanding
		   instructions for MAMA high voltage activation planned for 
		   this SMS.

		 - FSW and installation testing for MAMA operations is
		   on-going in the VEST/SEER today.  The FRR for this
		   upgrade is planned for next Tuesday, 22 April. 

	- SMS125 - Being worked this week.  The calendar is close to
		   completion.  It will contain the last piece of MAMA
		   HV Turn-on (7105) early in the week.

2.  Replans

	STIS

	- MAMA HV Turn-on (7105) is currently planned for SMS118.
	  Downstream MAMA scheduling is being assessed.  According to
	  the latest estimates, regular MAMA SMOV operations should start
	  in the SMS 139.

	NICMOS

	- No significant replans have yet been identified.


OPERATIONS STATUS

1.  OPUS - Problem in proper processing of STIS wavecals was reported and
	is being investigated.
  

	OPUS is requesting an independent check of their calculations for
	the "Re-use Target Offset" planned for a STIS ERO tommorrow morning.	
 	The STIS team will follow up.

2.  Archive - Currently down for a software installation V7.07.   Back up
	by noon.
 




biagetti@stsci.edu

Go Back Go back to the Meeting Minutes Page